# On the sample complexity of estimation in logistic regression

### v1

@@ 4 Discussion
Our characterization of the sample complexity of estimation in logistic regression deliniates the high, moderate, and low temperature regimes. However, we are only aware of computationally tractable estimators that achieve the (near) optimal sample complexities at high and zero temperatures: e.g., the linear estimator of Plan and Vershynin (2012) for $\beta\lesssim 1$, and the estimator based on solving a linear feasibility program (or the algorithm of Balcan and Long (2013)) for $\beta=\infty$. It would be interesting to determine if MLE achieves optimal sample complexity in the other regimes, and if the method of optimization is important—for instance, due to the implicit biases of the optimization procedure (Telgarsky, 2013; Soudry et al., 2018; Ji and Telgarsky, 2019).
@@ Appendix A Bounds on Gaussian integrals
$\Pr(\mathbf{z}\geq t)\leq\frac{1}{t\sqrt{2\pi}}\exp(-t^{2}/2).$
@@
$\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]\leq\min\left\{\frac{1}{2},\ \operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\right\}$
@@
$\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\leq\frac{1}{\beta}\sqrt{\frac{2}{\pi}}.$
@@
$g^{\prime\prime}(\eta)=\frac{1}{(1+\exp(\eta))(1+\exp(-\eta))}\leq\min\left\{\frac{1}{2},\ \exp(-\lvert\eta\rvert)\right\}.$
This establishes the first inequality. For the second inequality,
@@
$=2\exp\lparen\beta^{2}/2\rparen\Pr(\mathbf{z}\geq\beta)$ )
$\leq\frac{1}{\beta}\sqrt{\frac{2}{\pi}},$
where the inequality follows by Lemma 7. ∎
Remark 1.
Using a similar analysis as above, one can check that the bounds in Lemma 8 are fairly tight up to constants:
$\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]\geq\frac{1}{4}\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]$
and
[…]
@@
$\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}^{\prime})))\right]\leq\frac{\beta}{2}(1-\rho)\min\left\{\beta,2\sqrt{2/\pi}\right\}.$
@@
$\operatorname{\mathbb{E}}[g^{\prime}(\beta\mathbf{z})\beta\lparen\mathbf{z}-\mathbf{z}^{\prime}\rparen]$ $=\operatorname{\mathbb{E}}\left[g^{\prime}(\beta\mathbf{z})\beta\lparen 1-\rho\rparen\mathbf{z}-\sqrt{1-\rho^{2}}\mathbf{z}_{\perp}\right]$
@@
$=\beta^{2}\lparen 1-\rho\rparen\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]$
$\leq\frac{\beta}{2}\lparen 1-\rho\rparen\min\left\{\beta,2\sqrt{2/\pi}\right\},$
where the penultimate step follows by Stein’s identity, and the final step follows by Lemma 8. ∎
Lemma 11.
$\Pr(\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}<0)=\Pr(\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}>0)=\frac{\arccos(\rho)}{\pi}.$
@@
The claim is trivially true if $\lvert\rho\rvert=1$ since $\arccos(1)/\pi=0$ and $\arccos(-1)/\pi=1$. So without loss of generality assume $\lvert\rho\rvert<1$. Since $\mathbf{z}$ and $\mathbf{z}^{\prime}$ have correlation $\rho$, we may write
$\mathbf{z}^{\prime}=\rho\mathbf{z}+\sqrt{1-\rho^{2}}\mathbf{z}_{\perp},$
where $\mathbf{z}_{\perp}\sim\operatorname{N}(0,1)$ is independent of $\mathbf{z}$. Then, with $m=\rho/\sqrt{1-\rho^{2}}$,
$\Pr\lparen\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}<0\rparen=\Pr\lparen\mathbf{z}_{\perp}>-m\mathbf{z}\mid\mathbf{z}<0\rparen.$ ) .
Because the joint distribution of $(\mathbf{z},\mathbf{z}_{\perp})$ is spherically symmetric, the probability in the previous display is equal to the proportion of the “left half” ($x<0$) of the unit circle in the $x$-$y$ plane that lies above the $y=-mx$ line: this proportion is $\arccos(\rho)/\pi$. An analogous argument proves the same identity for $\Pr(\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}>0)$. ∎

### v3

@@ 4 Discussion
Our characterization of the sample complexity of estimation in logistic regression delineates the high, moderate, and low temperature regimes. However, we are only aware of computationally efficient estimators that achieve the (near) optimal sample complexities at high and zero temperatures: e.g., the linear estimator of Plan and Vershynin (2012) for $\beta\lesssim 1$, and the estimator based on solving a linear feasibility program (or the algorithm of Balcan and Long (2013)) for $\beta=\infty$. Note that, although the ReLU loss is convex, we need to minimize it over the sphere. It would be interesting to determine if the MLE itself (i.e., minimizing the logistic loss), or its efficient approximations, can shown to achieve optimal sample complexity.
Acknowledgements
We are most grateful to Akshay Krishnamurthy for helpful discussions about non-asymptotic analysis of MLE, and to the anonymous COLT reviewers for their constructive feedback and suggestions. Part of this work was done while DH was visiting the Halıcıoğlu Data Science Institute at UC San Diego in Spring 2023. We acknowledge the support of the National Science Foundation under grants IIS 2040971 and CCF 2217058/2133484.
@@ Appendix A Bounds on Gaussian integrals
($\left\lparen 1-\frac{1}{t^{2}}\right\rparen\frac{1}{t\sqrt{2\pi}}\exp(-t^{2}/2)\leq\Pr(\mathbf{z}\geq t)\leq\frac{1}{t\sqrt{2\pi}}\exp(-t^{2}/2).$ ) .
@@
$\frac{1}{4}\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\leq\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]\leq\min\left\{\frac{1}{2},\ \operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\right\}$ (3)
@@
[…]
@@
$g^{\prime\prime}(\eta)=\frac{1}{(1+\exp(\eta))(1+\exp(-\eta))}=\frac{1}{(1+\exp(\lvert\eta\rvert))(1+\exp(-\lvert\eta\rvert))}\leq\min\left\{\frac{1}{2},\ \exp(-\lvert\eta\rvert)\right\},$
while
$\frac{\exp(\lvert\eta\rvert)}{(1+\exp(\eta))(1+\exp(-\eta))}=\left\lparen\frac{1}{1+\exp(-\lvert\eta\rvert)}\right\rparen^{2}\geq\frac{1}{4}.$
Plugging in $\eta=\beta\mathbf{z}$ and taking expectations gives the upper- and lower-bounds on $\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]$. For the upper- and lower-bounds on $\operatorname{\mathbb{E}}[\exp(-\beta\lvert\mathbf{z}\rvert)]$, we have
@@
$=2\exp\lparen\beta^{2}/2\rparen\Pr(\mathbf{z}\geq\beta),$
and therefore the conclusion follows by applying Lemma 11. ∎
@@
Lemma 14.
Let $\mathbf{z}\sim\operatorname{N}(0,1)$. For any $\beta>0$ and any non-negative integer $q$,
$\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen\lvert\mathbf{z}\rvert^{q}]\leq\frac{q!}{\beta^{q}}.$
We remark that the bound in Lemma 14 can be improved to $O(q!/\beta^{q+1})$, which is tight up to constants. This is because $\operatorname{\mathbb{E}}[\exp(-\beta\lvert\mathbf{z}\rvert)\lvert\mathbf{z}\rvert^{q}]=\sqrt{2/\pi}\mu_{q}m_{\beta}$, where $\mu_{q}$ is the $q$-th uncentered moment of the $[0,\infty\rparen$ )-truncated $\operatorname{N}(-\beta,1)$ distribution, and $m_{\beta}=(1-\Phi(\beta))/\phi(\beta)$ is the Mills ratio for $\operatorname{N}(0,1)$ at $\beta$. However, we do not need this improved bound in the present work.
Proof of Lemma 14.
The function $f\colon\mathbb{R}_{+}\to\mathbb{R}$ defined by $f(x)=-\beta x+q\log(x)$ is maximized at $x=q/\beta$. Therefore
$\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen\lvert\mathbf{z}\rvert^{q}]$ $=\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert+q\ln\lvert\mathbf{z}\rvert\rparen]$
$\leq\exp\lparen-\beta(q/\beta)+q\ln(q/\beta)\rparen$ )
$=\exp\lparen-q+q\ln(q)-q\ln(\beta)\rparen$ )
$\leq\exp\lparen\ln(q!)-q\ln(\beta)\rparen$ )
$=\frac{q!}{\beta^{q}}.\qed$
@@
$\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}^{\prime})))\right]\leq\frac{\beta}{2}(1-\rho)\min\left\{\beta,2\sqrt{2/\pi}\right\}$
and
$\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}^{\prime})))\right]\geq\frac{\beta}{4}(1-\rho)\left\lparen 1-\frac{1}{\beta^{2}}\right\rparen\sqrt{\frac{2}{\pi}}.$
@@
[…]
@@
$=\beta^{2}\lparen 1-\rho\rparen\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})],$
where the final step follows by Stein’s identity. Now apply Lemma 12 to obtain the conclusion. ∎
Appendix B Proof of Lemma 5
The following lemma is implicit in the proofs of Lemma 5.2.1 and Lemma A.2.1 of Kuchelmeister and van de Geer (2023).
Lemma 16 (Kuchelmeister and van de Geer, 2023).
Fix $\theta^{\star},\theta\in S^{d-1}$ and $p\colon\mathbb{R}\to[0,1]$ satisfying $p(-t)=1-p(t)$ for all $t\in\mathbb{R}$. Let $\mathbf{x}$ be a standard normal random vector in $\mathbb{R}^{d}$; let the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ be $\operatorname{Bern}(p(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}))$; and define
$\boldsymbol{\delta}_{\theta}:=[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+}-[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}]_{+}.$
For any integer $q\geq 2$,
$\operatorname{\mathbb{E}}\left[\lvert\boldsymbol{\delta}_{\theta}\rvert^{q}\right]\leq 2^{q-1}\frac{1}{\pi\sqrt{2}}\Gamma\left\lparen\frac{q+1}{2}\right\rparen\left\lparen\frac{\pi}{\sqrt{2}}\lVert\theta-\theta^{\star}\rVert\right\rparen^{q+1}+2^{q-2}\lVert\theta-\theta^{\star}\rVert^{2q}\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\lvert\mathbf{z}\rvert^{q}\right]\\ +2^{2(q-1)}\lVert\theta-\theta^{\star}\rVert^{q}\left\lparen 1-\frac{1}{4}\lVert\theta-\theta^{\star}\rVert^{2}\right\rparen^{q/2}\frac{2^{q/2}}{\sqrt{\pi}}\Gamma\left\lparen\frac{q+1}{2}\right\rparen\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\right].$
@@ Appendix B Proof of Lemma 5
Using the identity $[x]_{+}=(\lvert x\rvert+x)/2$, we have for any $\theta\in S^{d-1}$,
$\boldsymbol{\delta}_{\theta}=[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+}-[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}]_{+}$ $=\frac{\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta\rvert-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta}{2}-\frac{\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}}{2}$
$=\frac{(\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta)-\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}))\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta}{2}-\frac{(\mathbf{y}-\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}))\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}(\theta-\theta^{\star})}{2}.$
Therefore, for any $q\geq 2$, Jensen’s inequality implies
$\lvert\boldsymbol{\delta}_{\theta}\rvert^{q}$ $\leq 2^{q-1}\left\lparen\left\lvert\frac{(\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta)-\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}))\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta}{2}\right\rvert^{q}+\left\lvert\frac{(\mathbf{y}-\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}))\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}(\theta-\theta^{\star})}{2}\right\rvert^{q}\right\rparen$ )
$=2^{q-1}\mathds{1}\{\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta)\neq\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\}\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta\rvert^{q}+2^{q-1}\mathds{1}\{\mathbf{y}\neq\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\}\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}(\theta-\theta^{\star})\rvert^{q}.$
Kuchelmeister and van de Geer (2023, Corollary A.2.1) showed that
[…]
where the latter inequality uses $\Gamma(q/2+1)/(q+1)\leq\Gamma((q+1)/2)$.
Using the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$,
$\operatorname{\mathbb{E}}\left[\mathds{1}\{\mathbf{y}\neq\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\}\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}(\theta-\theta^{\star})\rvert^{q}\right]$ $=\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}(\theta-\theta^{\star})\rvert^{q}\right]$
$=\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\lvert(\rho-1)\mathbf{z}+\sqrt{1-\rho^{2}}\mathbf{z}_{\perp}\rvert^{q}\right]$
where $\rho=\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$, and $\mathbf{z}$ and $\mathbf{z}_{\perp}$ are independent standard normal random variables. By Jensen’s inequality,
$\lvert(\rho-1)\mathbf{z}+\sqrt{1-\rho^{2}}\mathbf{z}_{\perp}\rvert^{q}\leq 2^{q-1}\lvert\rho-1\rvert^{q}\lvert\mathbf{z}\rvert^{q}+2^{q-1}(1-\rho^{2})^{q/2}\lvert\mathbf{z}_{\perp}\rvert^{q}.$
Moreover, we have
$1-\rho=\frac{1}{2}\lVert\theta-\theta^{\star}\rVert^{2}\quad\text{and}\quad 1-\rho^{2}=\lVert\theta-\theta^{\star}\rVert^{2}\left\lparen 1-\frac{1}{4}\lVert\theta-\theta^{\star}\rVert^{2}\right\rparen.$ ) .
Therefore, using independence of $\mathbf{z}$ and $\mathbf{z}_{\perp}$,
$\operatorname{\mathbb{E}}\left[\mathds{1}\{\mathbf{y}\neq\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\}\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}(\theta-\theta^{\star})\rvert^{q}\right]$
$\leq 2^{q-1}\lvert\rho-1\rvert^{q}\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\lvert\mathbf{z}\rvert^{q}\right]+2^{q-1}(1-\rho^{2})^{q/2}\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\right]\operatorname{\mathbb{E}}\left[\lvert\mathbf{z}_{\perp}\rvert^{q}\right]$
$=\frac{1}{2}\lVert\theta-\theta^{\star}\rVert^{2q}\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\lvert\mathbf{z}\rvert^{q}\right]+2^{q-1}\lVert\theta-\theta^{\star}\rVert^{q}\left\lparen 1-\frac{1}{4}\lVert\theta-\theta^{\star}\rVert^{2}\right\rparen^{q/2}\frac{2^{q/2}}{\sqrt{\pi}}\Gamma\left\lparen\frac{q+1}{2}\right\rparen\operatorname{\mathbb{E}}\left[p(-\lvert\mathbf{z}\rvert)\right].\qed$
Proof of Lemma 5.
We use Lemma 16 with $p(t)=g^{\prime}(\beta t)$. Therefore we need to bound $\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{z}\rvert)]$ and $\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{z}\rvert)\lvert\mathbf{z}\rvert^{q}]$ for all integers $q\geq 2$. Since $g^{\prime}(-\beta\lvert z\rvert)\leq\exp(-\beta\lvert z\rvert)$ for all $z\in\mathbb{R}$, we can use Lemma 12 for the former and Lemma 14 for the latter. We obtain
[…]
$\qquad+2^{q-2}\cdot\lVert\theta-\theta^{\star}\rVert^{2q}\cdot\frac{q!}{\beta^{q}}$
$\qquad+2^{2(q-1)}\cdot\lVert\theta-\theta^{\star}\rVert^{q}\left\lparen 1-\frac{1}{4}\lVert\theta-\theta^{\star}\rVert^{2}\right\rparen^{q/2}\frac{2^{q/2}}{\sqrt{\pi}}\Gamma\left\lparen\frac{q+1}{2}\right\rparen\cdot\frac{1}{\beta}\sqrt{\frac{2}{\pi}}$
[…]
$\qquad+2^{2(q-1)}\cdot\lVert\theta-\theta^{\star}\rVert^{q}\frac{2^{q/2}}{\sqrt{\pi}}\Gamma\left\lparen\frac{q+1}{2}\right\rparen\cdot\frac{1}{\beta}\sqrt{\frac{2}{\pi}}$
$\leq q!\cdot\left\lparen C\lVert\theta-\theta^{\star}\rVert\right\rparen^{q-2}\cdot\frac{C}{2}\cdot\lVert\theta-\theta^{\star}\rVert^{2}\cdot\left\lparen\lVert\theta-\theta^{\star}\rVert+\frac{1}{\beta}\right\rparen$ )
for some absolute constant $C>0$, where the final inequality uses $\Gamma((q+1)/2)\leq q!$. The final right-hand side in the previous display is clearly bounded above by $q!vb^{q-2}/2$ for the specified choices of $b$ and $v$. ∎
Appendix C Proof of Lemma 10
Proof of Lemma 10.
We have
$\operatorname{err}_{\theta^{\star}}(\theta)$ $=\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta)\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\right]$
$=\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta)\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\}\right]+\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta)\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}>0\}\right]$
$=\operatorname{\mathbb{E}}\left[(1-g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert))\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\}\right]+\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}>0\}\right]$
$=\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0)-2\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\}\right]+\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\right]$
$=\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0)-2\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\}\right]+\operatorname{err}_{\theta^{\star}}(\theta^{\star}).$
Therefore
$\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0)$ $=\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})+2\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\mathds{1}\{\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\}\right]$
$\leq\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})+2\operatorname{\mathbb{E}}\left[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)\right].$
The claim now follows by Lemma 13 and the fact that
$\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0)=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\geq\frac{\sqrt{2(1-\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}}{\pi}=\frac{\lVert\theta-\theta^{\star}\rVert}{\pi}.\qed$